# A Necessary Condition for the Submergence of Proto-Neutron Star Magnetic Fields by Supernova Fallback

### v1

@@ 5.1 Implications for the Diversity in Young isolated NS
We discuss how the crystallization of the fallback material affects Figure 7. We have derived Equations (24) and (25) by comparing $t_{\rm stall}^{\prime}$ with $t_{\rm fb}$. However, the submergence of the PNS field requires the new crust formation (see Figure 1), which means that we need to take into account the time interval from when the shock stalls to when the gas density and temperature satisfy the crystallizing condition. To estimate such a time interval, we have performed an additional simulation based on model BnMm2R10 (see Appendix A for details). From this simulation, we find that the time interval is of the same order as the shock stalling time scale. Therefore, we conclude that the crystallization of the fallback material does not significantly affect Figure 7. The crystallization time scale may become longer if we work with a more complete and realistic equation of state, as the degenerate pressure suppresses the increase in the gas density. Such simulations are left for important future work (Bernal et al., 2010, 2013).
5.2 Estimation of the effects of neutrino scattering and absorption
Our numerical models assume that neutrino cooling operates in an optically thin regime (see Equation 14). To evaluate the validity of this approximation, we estimate the radii of the neutrinospheres for $\nu_{\rm e}$ and $\bar{\nu}_{\rm e}$, at which the neutrino depths $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ become unity, respectively.
We consider neutrino scattering and absorption by free protons and neutrons. We ignore the effect of neutrino-electron scattering because its cross section is much smaller than those of neutrino-proton and neutrino-neutron scatterings (Shapiro & Teukolsky, 1983). We write the electron, proton, and neutron masses as $m_{\rm e}$, $m_{\rm p}$, and $m_{\rm n}$, respectively ($m_{\rm p}\approx m_{\rm n}\gg m_{\rm e}$). We define the cross sections for the absorption processes of $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ and $\nu_{e}+n\rightarrow e^{-}+p$ as
$\sigma_{\rm abs}(\bar{\nu}_{\rm e}p)$ $=$ $\sigma_{0}\left(\frac{1+3g_{\rm A}^{2}}{4}\right)\left(\frac{\varepsilon_{\bar{\nu}_{\rm e}}-\Delta_{\rm np}}{m_{\rm e}}\right)^{2}\left[1-\left(\frac{m_{\rm e}}{\varepsilon_{\bar{\nu}_{\rm e}}-\Delta_{\rm np}}\right)^{2}\right]^{1/2},$ (28)
$\sigma_{\rm abs}(\nu_{\rm e}n)$ $=$ $\sigma_{0}\left(\frac{1+3g_{\rm A}^{2}}{4}\right)\left(\frac{\varepsilon_{\nu_{\rm e}}+\Delta_{\rm np}}{m_{\rm e}}\right)^{2}\left[1-\left(\frac{m_{\rm e}}{\varepsilon_{\nu_{\rm e}}+\Delta_{\rm np}}\right)^{2}\right]^{1/2},$ (29)
respectively (Burrows et al., 2006). Here, $\sigma_{0}=1.705\times 10^{-44}~{}{\rm cm^{2}}$, $g_{\rm A}$ is the axial–vector coupling constant ($g_{\rm A}\approx-1.23$), $\varepsilon_{\nu_{\rm e}}$ is the energy of $\nu_{\rm e}$, $\varepsilon_{\bar{\nu}_{\rm e}}$ is the energy of $\bar{\nu}_{\rm e}$, and $\Delta_{\rm np}=m_{\rm n}-m_{\rm p}$. We assume that the correction for weak magnetism and recoil is negligible for simplicity. We describe the cross sections in the elastic limit of neutrino-proton and neutrino-neutron scatterings as follows (see e.g., Shapiro & Teukolsky, 1983; Suwa et al., 2019):
$\sigma_{\rm sca,p}\sim\sigma_{\rm sca,n}=\frac{\sigma_{0}}{4}\left(\frac{\varepsilon_{\nu}}{m_{\rm e}}\right)^{2}.$ (30)
The method for estimating the radii of the neutrinospheres is as follows. We denote the number densities of electrons, protons, and neutrons as $n_{\rm e}$, $n_{\rm p}$, and $n_{\rm n}$, respectively. We assume the charge neutrality condition $n_{\rm e}=n_{\rm p}$. We can then write the mass density as $\rho=m_{\rm p}(n_{\rm p}+n_{\rm n})$. We represent the electron fraction as $Y_{\rm e}=n_{\rm p}/(n_{\rm p}+n_{\rm n})$. From these definitions, we get the following relations: $n_{\rm p}=\rho Y_{\rm e}/{m_{\rm p}}$ and $n_{\rm n}={(1-Y_{\rm e})}\rho/m_{\rm p}$. The detailed estimation of the value of $Y_{e}$ is beyond the scope of this study, and we adopt $Y_{\rm e}=0.5$ as a fiducial value for simplicity (Fryer et al., 2006). From the definitions explained above, we calculate the radii of neutrinospheres using the following equations:
[…]
Since the typical gas temperature at $r\sim r_{\rm M}$ is $T\sim 10^{11}~{}{\rm K}$, we calculate $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ for $\varepsilon_{\nu_{\rm e}}=\varepsilon_{\bar{\nu}_{\rm e}}=10~{}{\rm MeV}$ (see also Akaho et al., 2024).
In all models, when $t=t_{\rm stall}$, both $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ are much smaller than unity even at $r=r_{\rm M}$, suggesting that neutrino scattering and absorption by free protons and neutrons can be neglected (see Sumiyoshi & Yamada, 2012, for more detailed estimation).
@@ 6 Summary
We developed semi-analytic models that are consistent with the magnetospheric and shock radii ($r_{\rm M}$ and $r_{\rm sh}$, respectively) when the shock stalls (Section 4.2). The semi-analytic models are based on the following two relations: (1) the pressure balance between the magnetic pressure of the PNS field and the radiation pressure in the post-shock region (Equation 16), and (2) the balance between the release rate of the gravitational energy and energy loss rate due to neutrino cooling at $r=r_{\rm M}$ (Equation 17). Therefore, both $r_{\rm M}$ and $r_{\rm sh}$ decrease with $\dot{M}_{\rm fb}$ and increase with $B_{\rm PNS}$ (see top and middle panels of Figure 5). Our semi-analytic model reduces to the analytical solution in Chevalier (1989) when $B_{\rm PNS}=0$.
@@
Based on the necessary condition, we updated possible solutions of the supernova fallback to better understand the origin of the diversity of NS magnetic fields (Figure 7 in Section 5.1). The solutions require further updates because they ignore some multidimensional effects. However, we believe that the presented scheme will be a new benchmark for the scenarios of the NS diversity.
@@
We thank Ryuichiro Akaho, Yudai Suwa, and Kenta Dohi for fruitful discussions. This work was supported by JSPS KAKENHI grant Nos. JP24KJ0143, JP25K17439 (A.I.), JP22K14074, JP22KK0043, JP21H04487 (S.T.), JP24K00668, JP23H04899, JP22H00130 (K.K.), and JP24K00672, JP21H04488, JP24K00678 (H.R.T). A part of this research has been funded by the MEXT as ”Program for Promoting Researches on the Supercomputer Fugaku” (Toward a unified view of the universe: from large-scale structures to planets, JPMXP1020200109; A.I. and H.R.T.). Numerical computations were performed with computational resources provided by Cray XC 50 at the Center for Computational Astrophysics (CfCA) of the National Astronomical Observatory of Japan (NAOJ), the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, The University of Tokyo.
Appendix A Time scale for the crust formation
@@
When Coulomb energy dominates the thermal energy in the accreted matter, ions will crystallize into a periodic lattice structure to form the NS crust. The crystallization of ions occurs if the gas temperature cools down to a melting temperature $T_{\rm m}=(Z^{2}e_{\rm c}^{2})(\Gamma_{\rm m}k)^{-1}(4\pi n_{\rm ion}/3)^{1/3}$ (see e.g., Shapiro & Teukolsky, 1983; Chamel & Haensel, 2008), where $Z$ is the typical proton number of the lattice component, $e_{\rm c}$ is the elementary charge, $\Gamma_{\rm m}$ is the ratio of Coulomb to thermal energies of the lattice at the melting point, $k$ is the Boltzmann constant, and $n_{\rm ion}$ is the number density of ion. For simplicity, we assume that the crust is formed of a perfect crystal with a single nuclear species at lattice sites (Chamel & Haensel, 2008). Its mass number and mass fraction are denoted by $A=Z/Y_{\rm e}$ and $x_{\rm a}$, respectively. We then write the number density of ions as $n_{\rm ion}=(\rho x_{\rm a})/(Am_{\rm p})$. From these definitions, we can rewrite the melting temperature as (Suwa, 2014):
$T_{\rm m}=5.6\times 10^{9}~{}[{\rm K}]\left(\frac{\Gamma_{\rm m}}{175}\right)^{-1}\left(\frac{\rho}{10^{14}~{}{\rm g~{}cm^{-3}}}\right)^{1/3}\left(\frac{Y_{\rm e}}{0.5}\right)^{1/3}\left(\frac{x_{\rm a}}{0.1}\right)^{1/3}\left(\frac{Z}{28}\right)^{5/3}.$ (A1)
We adopt $\Gamma_{\rm m}=175$ and $x_{\rm a}=0.1$ as fiducial values (Suwa, 2014). We consider nuclei with the magic neutron numbers $Z=28$, $50$, and $82$ (Haensel, 2001; Chamel & Haensel, 2008).
To estimate the time scale for crystallization, we continue the simulation in model BnMm2R10 until the temperature reaches the melting temperature defined by Equation (A1). The details of the setup are as follows. To ensure sufficient spatial resolution while maintaining reasonable computational cost, we conduct the simulation in three successive stages. We first set $r_{\rm out}$ to $20~{}{\rm km}$ with the number of numerical grid points of $N_{r}=16384$. Physical quantities at each grid point are initialized by linearly interpolating the data from model BnMm2 at time $t\approx 0.0354~{}{\rm s}$. For grid points located inside the innermost radial point of model BnMm2, we assign the values of the physical quantities at the innermost point of model BnMm2. We run a simulation by $t=0.0356~{}{\rm s}$, and then the resulting profile is used as the initial condition for the next stage. This procedure is repeated for $r_{\rm out}=10.1~{}{\rm km}$ and $10.01~{}{\rm km}$. This approach ensures $\Delta r<H/10$.
Figure 8: Left: Radial profiles of the gas density (top) and the gas temperature (bottom) in the vicinity of the PNS surface. Right: Plots for the gas temperature against the gas density (black line). Crystallizing temperature in the cases of $Z=28$ (solid), $Z=50$ (dotted), $Z=82$ (dashed) are represented by red lines (see Equation A1). All plots are measured at $t\approx 0.0357~{}{\rm s}$.

### v2

@@ 5.1 Implications for the Diversity in Young isolated NS
We discuss how the crystallization of the fallback material affects Figure 7. We have derived Equations (24) and (25) by comparing $t_{\rm stall}^{\prime}$ with $t_{\rm fb}$. However, the submergence of the PNS field requires the new crust formation (see Figure 1), which means that we need to take into account the time interval from when the shock stalls to when the gas density and temperature satisfy the crystallizing condition. To estimate such a time interval, we have performed an additional simulation based on model BnMm2R10 (see Appendix B for details). The simulation shows that the time interval is of the same order as the shock stalling time scale. Therefore, we conclude that the crystallization of the fallback material does not significantly affect Figure 7. The crystallization time scale may become longer if we work with a more complete and realistic EoS (Suwa2014; Nakazato2025), as the degenerate pressure suppresses the increase in the gas density. We will discuss this point in Section 5.2.
5.2 Model Limitations
Regarding Figure 7, we discuss possible multidimensional effects, which are ignored in our simulations. Recent multidimensional simulations of the supernova explosion suggested that the fallback gas may have a large angular momentum (see Janka2022, and references therein). When such material accretes onto the PNS, an accretion disk will form around the PNS magnetosphere (Takahashi2017; Inoue2023; Inoue2024). The differential rotation of the disk opens up the PNS dipole field, which enhances $L_{\rm spin}$ (Parfrey2017). Furthermore, MHD instabilities may also play an important role. The surface at $r_{\rm M}$, where the magnetic pressure balances the radiation pressure, can be unstable to the magnetic Rayleigh-Taylor instability (see e.g., Cumming2001; Stone2007). This instability leads to dynamic mixing around $r\sim r_{\rm M}$, suppressing the submergence of the PNS magnetosphere (Bernal2013). Therefore, the $\dot{M}_{\rm fb}$ required for the submergence will increase.
Realistic geometries of the dipole magnetic field and accretion flows can also lead to multidimensional effects. The large-scale field geometry and magnetic tension, which are ignored in the present study, can form accretion mountains around the magnetic poles (Melatos2001; Payne2004; Payne2007; Mukherjee2017). As the mountains spread, they will compress the PNS magnetic fields toward the equator (Fujisawa2022; Yeole2025), thereby enhancing the local magnetic pressure and suppressing the submergence in this region. Furthermore, multidimensional simulations of core-collapse supernova showed that neutrino-driven wind and accretion occur at the same time (Bollig2021; Nakamura2025; Vartanyan2025), implying the presence of anisotropic fallback accretion flows. Such anisotropic flows may partially submerge the dipole field, forming an NS with complex small-scale multipole fields (Payne2004; Suvorov2020). When the NS spins down via multipole radiation, its spin-down rate is smaller than that of dipole radiation at a fixed magnetic field strength. Therefore, the NS will be identified as the CCO if its spin-down is interpreted as originating from the dipole radiation. The partial submergence can be achieved with a fallback rate smaller than that derived in this study.
The long-term evolution of the accreted material remains unexplored due to the limited duration of our simulations. This study has focused on the submergence during the PNS phase, in which the remnant formed by core collapse (i.e., the PNS) contracts via neutrino cooling. This phase lasts for several tens of seconds after the core bounce. However, a submergence may also occur in a stage much later than the PNS phase. The post-shock gas redistributes itself to maintain hydrostatic equilibrium after the fallback phase for $t_{\rm stall}^{\prime}>t_{\rm fb}$ (see, Section 4.5). The hydrostatic atmosphere will then cool via neutrino emission (Combi2025) and/or radiation (Takahashi2018; Abarca2018) on a time scale much longer than the duration of the PNS phase. The atmosphere may submerge the magnetic fields at the crustal level in the post-PNS phase. The necessary condition derived in this study does not account for such late-time submergence. Therefore, we cannot rule out the possibility that the late-time submergence may occur even if the necessary condition is not satisfied. To clarify the scenario, simulations that trace the long-term evolution of the atmosphere are required.
Finally, we make a few remarks about caveats for the EoS adopted in this study. We assume a specific heat ratio of $\Gamma=4/3$ for the EoS and calculate the temperature from $T=(3p/a_{\rm rad})^{1/4}$ (Chevalier1989). However, this is still an approximation. To assess the validity of this approximation, we perform an additional simulation for $r_{\rm NS}=10~{\rm km}$ using the Helmholtz EoS (Timmes2000) (see Appendix C for details). The simulation shows that the pair pressure increases the fluid pressure by up to a factor of two, resulting in an increase in the shock radius by a similar factor (Houck1991). The resulting shock stalling time scale is longer than that in the original case with $r_{\rm NS}=10~{\rm km}$, but shorter than that in the original case with $r_{\rm NS}=20~{\rm km}$. Therefore, the realistic $t_{\rm stall}^{\prime}=t_{\rm fb}$ line for $r_{\rm NS}=10~{\rm km}$ will lie between those for $r_{\rm NS}=10~{\rm km}$ and for $r_{\rm NS}=20~{\rm km}$ in Figure 7. In addition, the additional simulation shows that while the electron degeneracy pressure is dynamically negligible for $t\leq t_{\rm stall}$, it becomes non-negligible for $t>t_{\rm stall}$ (Bernal2010). Therefore, the pressure will increase the crystallization time scale. Since the qualitative behavior and overall trends remain unchanged, we retain the ideal-gas approximation for the remaining parameter survey. To summarize this section, the solutions in Figure 7 require further updates. However, we believe that our results will offer valuable guidance in investigating the diversity of young isolated NSs through multidimensional simulations.
@@ 6 Summary
We developed semi-analytic models that are consistent with the magnetospheric and shock radii ($r_{\rm M}$ and $r_{\rm sh}$, respectively) when the shock stalls (Section 4.2). The semi-analytic models are based on the following two relations: (1) the pressure balance between the magnetic pressure of the PNS field and the radiation pressure in the post-shock region (Equation 16), and (2) the balance between the release rate of the gravitational energy and energy loss rate due to neutrino cooling at $r=r_{\rm M}$ (Equation 17). Therefore, both $r_{\rm M}$ and $r_{\rm sh}$ decrease with $\dot{M}_{\rm fb}$ and increase with $B_{\rm PNS}$ (see top and middle panels of Figure 5). Our semi-analytic model reduces to the analytical solution in Chevalier1989 when $B_{\rm PNS}=0$.
@@
Based on the necessary condition, we updated possible solutions of the supernova fallback to better understand the origin of the diversity of NS magnetic fields (Figure 7 in Section 5.1). The solutions require further updates because they ignore some multidimensional effects and the electron-positron pair pressure. However, we believe that the presented scheme will be a new benchmark for the scenarios of the NS diversity.
@@
We thank Ryuichiro Akaho, Yudai Suwa, and Akira Dohi for fruitful discussions. This work was supported by JSPS KAKENHI grant Nos. JP24KJ0143, JP25K17439 (A.I.), JP22K14074, JP22KK0043, JP21H04487 (S.T.), JP24K00668, JP23H04899, JP22H00130 ($K.$), and JP24K00672, JP21H04488, JP24K00678 (H.R.T). A part of this research has been funded by the MEXT as ”Program for Promoting Researches on the Supercomputer Fugaku” (Toward a unified view of the universe: from large-scale structures to planets, JPMXP1020200109; A.I. and H.R.T.). Numerical computations were performed with computational resources provided by Cray XC 50 at the Center for Computational Astrophysics (CfCA) of the National Astronomical Observatory of Japan (NAOJ), the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, The University of Tokyo.
Appendix A Estimation of the effects of neutrino scattering and absorption
Our numerical models assumed that neutrino cooling operates in an optically thin regime (see Equation 14). To evaluate the validity of this approximation, we estimate the radii of the neutrinospheres for $\nu_{\rm e}$ and $\bar{\nu}_{\rm e}$, at which the neutrino depths $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ become unity, respectively.
We consider neutrino scattering and absorption by free protons and neutrons. We neglect the effect of neutrino-electron scattering because its cross section is much smaller than those of neutrino-proton and neutrino-neutron scatterings (Shapiro1983). We write the electron, proton, and neutron masses as $m_{\rm e}$, $m_{\rm p}$, and $m_{\rm n}$, respectively ($m_{\rm p}\approx m_{\rm n}\gg m_{\rm e}$). We define the cross sections for the absorption processes of $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ and $\nu_{e}+n\rightarrow e^{-}+p$ as
$\sigma_{\rm abs}(\bar{\nu}_{\rm e}p)$ $=$ $\sigma_{0}\left(\frac{1+3g_{\rm A}^{2}}{4}\right)\left(\frac{\varepsilon_{\bar{\nu}_{\rm e}}-\Delta_{\rm np}}{m_{\rm e}}\right)^{2}\left[1-\left(\frac{m_{\rm e}}{\varepsilon_{\bar{\nu}_{\rm e}}-\Delta_{\rm np}}\right)^{2}\right]^{1/2},$ (A1)
$\sigma_{\rm abs}(\nu_{\rm e}n)$ $=$ $\sigma_{0}\left(\frac{1+3g_{\rm A}^{2}}{4}\right)\left(\frac{\varepsilon_{\nu_{\rm e}}+\Delta_{\rm np}}{m_{\rm e}}\right)^{2}\left[1-\left(\frac{m_{\rm e}}{\varepsilon_{\nu_{\rm e}}+\Delta_{\rm np}}\right)^{2}\right]^{1/2},$ (A2)
respectively (Burrows2006). Here, $\sigma_{0}=1.705\times 10^{-44}~{\rm cm^{2}}$, $g_{\rm A}$ is the axial–vector coupling constant ($g_{\rm A}\approx-1.23$), $\varepsilon_{\nu_{\rm e}}$ is the energy of $\nu_{\rm e}$, $\varepsilon_{\bar{\nu}_{\rm e}}$ is the energy of $\bar{\nu}_{\rm e}$, and $\Delta_{\rm np}=m_{\rm n}-m_{\rm p}$. We assume that the correction for weak magnetism and recoil is negligible for simplicity. We describe the cross sections in the elastic limit of neutrino-proton and neutrino-neutron scatterings as follows (see e.g., Shapiro1983; Suwa2019):
$\sigma_{\rm sca,p}\sim\sigma_{\rm sca,n}=\frac{\sigma_{0}}{4}\left(\frac{\varepsilon_{\nu}}{m_{\rm e}}\right)^{2}.$ (A3)
The method for estimating the radii of the neutrinospheres is as follows. We denote the number densities of electrons, protons, and neutrons as $n_{\rm e}$, $n_{\rm p}$, and $n_{\rm n}$, respectively. We can then write the mass density as $\rho=m_{\rm p}(n_{\rm p}+n_{\rm n})$. We assume the charge neutrality condition $n_{\rm e}=n_{\rm p}$. We represent the electron fraction as $Y_{\rm e}=n_{\rm p}/(n_{\rm p}+n_{\rm n})$. From these definitions, we get the following relations: $n_{\rm p}=\rho Y_{\rm e}/{m_{\rm p}}$ and $n_{\rm n}={(1-Y_{\rm e})}\rho/m_{\rm p}$. The detailed estimation of the value of $Y_{e}$ is beyond the scope of this study, and we adopt $Y_{\rm e}=0.5$ as a fiducial value for simplicity (Fryer2006). From the definitions explained above, we calculate the radii of neutrinospheres using the following equations:
$\tau_{\nu_{\rm e}}(\varepsilon_{\nu_{\rm e}})$ $=$ $\int^{\infty}_{r}(\sigma_{\rm abs}(\nu_{\rm e}n){n_{\rm n}}+\sigma_{{\rm sca},{\rm p}}{n_{\rm p}}+\sigma_{{\rm sca},{\rm n}}{n_{\rm n}})dr,$ (A4)
$\tau_{\bar{\nu}_{\rm e}}(\varepsilon_{\bar{\nu}_{\rm e}})$ $=$ $\int^{\infty}_{r}(\sigma_{\rm abs}(\bar{\nu}_{\rm e}p){n_{\rm p}}+\sigma_{{\rm sca},{\rm p}}{n_{\rm p}}+\sigma_{{\rm sca},{\rm n}}{n_{\rm n}})dr.$ (A5)
Since the typical gas temperature at $r\sim r_{\rm M}$ is $T\sim 10^{11}~{\rm K}$, we calculate $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ for $\varepsilon_{\nu_{\rm e}}=\varepsilon_{\bar{\nu}_{\rm e}}=10~{\rm MeV}$ (see also Akaho2024).
In all models, when $t=t_{\rm stall}$, both $\tau_{\nu_{\rm e}}$ and $\tau_{\bar{\nu}_{\rm e}}$ are much smaller than unity even at $r=r_{\rm M}$, suggesting that neutrino scattering and absorption by free protons and neutrons can be neglected (see Sumiyoshi2012, for more detailed estimation).
Appendix B Time scale for the crust formation
@@
When Coulomb energy dominates the thermal energy in the accreted matter, ions will crystallize into a periodic lattice structure to form the NS crust. The crystallization of ions occurs if the gas temperature cools down to a melting temperature $T_{\rm m}=(Z^{2}e_{\rm c}^{2})(\Gamma_{\rm m}k)^{-1}(4\pi n_{\rm ion}/3)^{1/3}$ (see e.g., Shapiro1983; Chamel2008), where $Z$ is the typical proton number of the lattice component, $e_{\rm c}$ is the elementary charge, $\Gamma_{\rm m}$ is the ratio of Coulomb to thermal energies of the lattice at the melting point, $k$ is the Boltzmann constant, and $n_{\rm ion}$ is the number density of ion. For simplicity, we assume that the crust is formed of a perfect crystal with a single nuclear species at lattice sites (Chamel2008). Its mass number and mass fraction are denoted by $A=Z/Y_{\rm e}$ and $x_{\rm a}$, respectively. We then write the number density of ions as $n_{\rm ion}=(\rho x_{\rm a})/(Am_{\rm p})$. From these definitions, we can rewrite the melting temperature as (Suwa2014):
$T_{\rm m}=5.6\times 10^{9}~[{\rm K}]\left(\frac{\Gamma_{\rm m}}{175}\right)^{-1}\left(\frac{\rho}{10^{14}~{\rm g~cm^{-3}}}\right)^{1/3}\left(\frac{Y_{\rm e}}{0.5}\right)^{1/3}\left(\frac{x_{\rm a}}{0.1}\right)^{1/3}\left(\frac{Z}{28}\right)^{5/3}.$ (B1)
We adopt $\Gamma_{\rm m}=175$ and $x_{\rm a}=0.1$ as fiducial values (Suwa2014). We consider nuclei with the magic numbers $Z=28$, $50$, and $82$ (Haensel2001; Chamel2008).
To estimate the time scale for crystallization, we continue the simulation in model BnMm2R10 until the temperature reaches the melting temperature defined by Equation (B1). The details of the setup are as follows. To ensure sufficient spatial resolution while maintaining reasonable computational cost, we conduct the simulation in three successive stages. We first set $r_{\rm out}$ to $20~{\rm km}$ with the number of numerical grid points of $N_{r}=16384$. Physical quantities at each grid point are initialized by linearly interpolating the data from model BnMm2 at time $t\approx 0.0354~{\rm s}$. For grid points located inside the innermost radial point of model BnMm2, we assign the values of the physical quantities at the innermost point of model BnMm2. We run a simulation by $t=0.0356~{\rm s}$, and then the resulting profile is used as the initial condition for the next stage. This procedure is repeated for $r_{\rm out}=10.1~{\rm km}$ and $10.01~{\rm km}$. This approach ensures $\Delta r<H/10$.
Figure 8: Left: Radial profiles of the gas density (top) and the gas temperature (bottom) in the vicinity of the PNS surface. Right: Plots for the gas temperature against the gas density (black line). Crystallizing temperature in the cases of $Z=28$ (solid), $Z=50$ (dotted), $Z=82$ (dashed) are represented by red lines (see Equation B1). All plots are measured at $t\approx 0.0357~{\rm s}$.
@@
Appendix C Results considering Helmholtz Equation of State
We assumed an effective adiabatic index of $\Gamma=4/3$ for the EoS and calculated the temperature from the Stefan-Boltzmann law for simplicity. To assess the impact of the EoS on our conclusions, we perform an additional test simulation employing a more realistic EoS.
We adopt the Helmholtz EoS (Timmes2000) for the additional simulation, which is based on a tabulated interpolation of the Helmholtz free energy computed by the Timmes EoS (Timmes1999). The Helmholtz EoS includes contributions from blackbody radiation, fully ionized nuclei, and degenerate and relativistic electrons and positrons. The fluid pressure and its internal energy density are obtained by summing over these components. When the temperature and density are specified, the Helmholtz EoS package returns the corresponding fluid pressure and internal energy density.
We describe the numerical methods and setup below. The total pressure is used as a primitive variable in the default version of the UWABAMI code, whereas we instead use the temperature in this simulation. We adopt the recovery scheme for the primitive variables proposed by Noble2006 (see also Siegel2018). The fluid is assumed to consist of free protons, neutrons, and electrons, with $Y_{\rm e}=0.5$. The PNS radius is set to $r_{\rm PNS}=10~{\rm km}$. We investigate the case with $B_{\rm PNS}=0$ and $\dot{M}_{\rm fb}=10^{-3}~{\rm M}\odot~{\rm s}^{-1}$. We refer to this model as the Helmholtz model and compare it with models BnMm3R10 and BnMm3R20.
Figure 9: The radial profiles of the gas density (upper panel) and temperatures (lower panel) when the shock stalls. The thick lines represent the results in the Helmholtz model, while the thin solid and the dashed lines indicate the results in models BnMm3R10 and BnMm3R20, respectively. The blue lines in the lower panel indicate the Fermi temperature.
The upper panel of Figure 9 shows the radial profiles of the gas density when the shock stalls. The thick line represents the result of the Helmholtz model, while the thin solid line and the dashed line indicate the results in models BnMm3R10 and BnMm3R20, respectively. In the post shock region, the resulting gas density in the Helmholtz model is greater than that in model BnMm3R10.
The upper panel of Figure 9 demonstrates that the shock radius in the Helmholtz model is greater than that in the ideal case with $r_{\rm NS}=10~{\rm km}$, but smaller than that in the ideal case with $r_{\rm NS}=20~{\rm km}$. The pair pressure increases the fluid pressure by up to a factor of two, resulting in a larger shock radius (Houck1991). The resulting shock stalling time scale is longer than that in the original case with $r_{\rm NS}=10~{\rm km}$, but shorter than that in the original case with $r_{\rm NS}=20~{\rm km}$.
The black lines in the lower panel of Figure 9 represent the radial profiles of the gas temperature. The resulting temperature in the Helmholtz model is very close to that in the original case with $r_{\rm NS}=10~{\rm km}$ near the PNS surface.
We compare the resulting temperature with the Fermi temperature. The blue lines in the lower panel of Figure 9 indicate the Fermi temperature. We calculate the Fermi momentum from $p_{\rm F}=(3\pi^{2}n_{\rm e})^{1/3}\hbar$. The Fermi energy is given by $E_{\rm F}=\sqrt{(p_{\rm F}c)^{2}+(m_{\rm e}c^{2})^{2}}-m_{\rm e}c^{2}$. The Fermi temperature is described as $T_{\rm F}=E_{\rm F}/k$. We find $T>T_{\rm F}$ for $t\leq t_{\rm stall}$ but $T<T_{\rm F}$ for $t>t_{\rm stall}$. Therefore, the degenerate pressure of the electrons is negligible for $t\leq t_{\rm stall}$. We note that the degenerate pressure cannot be neglected after the shock stalls. Because the degenerate pressure suppresses the increase in the gas density, the crystallization time scale (see Appendix B) will increase when we consider the Helmholtz EoS. The relation $T<T_{\rm F}$ for $t>t_{\rm stall}$ is consistent with the result in Bernal2010.